# https://www.mmlab-ntu.com/project/edgesam

### v2

@@ Appendix 0.B More Ablation Studies
Training Losses for Decoder. In our main paper, we have mentioned that supervising the student with teacher mask output as the ground truth is the most effective for decoder loss. Here, we provide more details of our empirical findings. In particular, we compare various configurations, including (1) combining mask loss with aligning attention maps from the teacher and student decoder; (2) combining mask loss with aligning the IoU predictions between teacher and student; (3) aligning the output feature maps between the teacher and student decoder. As shown in Tab. 6(b), the plain mask loss supervision yields the best performance. Thus, instead of exploring different loss combinations, we study the strategy to select the proper prompts during distillation and propose prompt-in-the-loop distillation.
@@
Number of Prompts. Given that each sample of SA-1B contains around 100 instances, training all of them in a single batch causes VRAM overflow. Therefore, we randomly sample a subset for each iteration. In Tab. 6(c), we study the impact of the number of prompts for sampling and find that 16 prompts achieve the best trade-off. Thus, we set it as the default.
Appendix 0.C Qualitative Results Compared with SAM
@@
Appendix 0.D Algorithm of Prompt-In-the-Loop Distillation
In Algo. 1, we provide the algorithmic pseudocode for the proposed prompt-in-the-loop distillation.
$T_{enc},S_{enc}\leftarrow\text{SAM / EdgeSAM encoder}$;
$T_{dec}\leftarrow\text{SAM decoder}$;
$S_{dec}\leftarrow\text{EdgeSAM decoder initialized w/}\ T_{dec}$;
$m,\mathbf{c}\leftarrow\text{shared mask / IoU tokens}$;
[…]
for $i=1,2,\dots,N$ do
       $f_{t},f_{s}\leftarrow T_{enc}(\mathcal{I}_{i}),S_{enc}(\mathcal{I}_{i})$;
       $p\leftarrow\text{select the box or point prompt in}\ \mathcal{P}_{i}$;
       $L\leftarrow L_{\text{mask}}(m_{t},m_{s})$;
       for $j=1,2,\dots,M$ do
             $\hat{p}\leftarrow\text{sample\_in\_disagree}(m_{t},m_{s})$;
             $p\leftarrow\hat{p}\text{ appends to}\ p$;
             $L\leftarrow L+L_{\text{mask}}(m_{t},m_{s})$;
       end for
      $S_{enc},S_{dec}\leftarrow\text{SGD model update}$;
end for
Algorithm 1 Prompt-In-the-Loop Distillation

### v3

@@ 4.3 Ablation Studies
Training Losses for Decoder. By default, we supervise the student with teacher mask output as the ground truth is the most effective for decoder loss. Here, we provide more details of our empirical findings. In particular, we compare various configurations, including (1) combining mask loss with aligning attention maps from the teacher and student decoder; (2) combining mask loss with aligning the IoU predictions between teacher and student; (3) aligning the output feature maps between the teacher and student decoder. As shown in Tab. 5(j), the plain mask loss supervision yields the best performance. Thus, instead of exploring different loss combinations, we study the strategy to select the proper prompts during distillation and propose prompt-in-the-loop distillation.
Train Stages. Due to the optimal optimization settings being different for encoder-only distillation and prompt-in-the-loop distillation, we find joint-training yields sub-optimal results compared to training in two stages (see Tab. 5(k)). Therefore, we propose to train them sequentially.
Training Datasets. In our default setting, we train EdgeSAM with 1% images of the SA-1B (Kirillov et al, 2023) dataset. In Tab. 5(l), we study the impact of training datasets on transferability. Specifically, we keep the model and training pipeline the same while switching the training dataset among 1% SA-1B, COCO (Lin et al, 2014), and LVIS (Gupta et al, 2019). Note that all datasets contain a similar number of images (approximately 110K). Results show that training with SA-1B archives the best performance across all test sets with box prompts, while with point prompts, training and testing in the same domain yield the best result. This suggests that with informative prompts, such as boxes, distillation on the SA-1B dataset is sufficient for zero-shot transfer. However, with ambiguous prompts, such as a single point, efforts must be made to capture the granularity priors of the testing datasets. This motivates us to leverage an additional RPN to capture such priors explicitly.
Figure 3: Qualitative results of EdgeSAM with point and box prompts. The green and red stars indicate the positive and negative points respectively.
@@
4.4 Qualitative Results
In Fig. 3, we visualize the mask quality of EdgeSAM with one point, two points, and one box as the prompts, respectively. Note that none of the images are from the SA-1B dataset that EdgeSAM is trained on, demonstrating its zero-shot generalization ability.
4.5 Effectiveness of the Prompt-in-the-Loop Knowledge Distillation
In Fig. 4, we compare EdgeSAM with MobileSAM with both points and boxes as prompts. Since EdgeSAM and MobileSAM adopt different backbones, we add a baseline that follows the same network architecture as EdgeSAM but is trained with MobileSAM’s encoder-only distillation method. Compared with task-agnostic KD methods, models trained with prompt-KD follow the user’s intent more precisely, e.g., MobileSAM and baseline fail to segment both persons on the bench. Besides, prompt-KD involves implicit hard example mining by iteratively sampling new prompts in the wrongly segmented area, which leads to tighter boundaries and more accurate results in difficult regions, e.g., the sculpture and bicycle in Fig. 4. These results align well with our assumption that task-agnostic encoder distillation fails to capture the full knowledge embodied in SAM and involving prompts in the loop during distillation helps capture the intricate dynamics between user input and mask generation.
4.6 Qualitative Results Compared with SAM
@@
5 Conclusion and Discussion
In this paper, we present EdgeSAM, the first SAM variant that runs in real-time on edge devices. We achieve this by distilling SAM into a lightweight CNN-based architecture. Our preliminary experiments show that existing distillation schemes, which only involve the image encoder, thus being task-agnostic, fail to reveal the full knowledge spectrum of SAM to the student model. Thus, we propose a prompt-in-the-loop knowledge distillation method that considers both the encoder and decoder of SAM and provides task-specific supervision signals. Experiments across SA-1B, COCO, and LVIS under various settings demonstrate the effectiveness of EdgeSAM in terms of efficiency and accuracy compared with SAM and other accelerated variants.
Despite being able to run EdgeSAM at a decent speed on edge devices, we still have several research directions, that could potentially provide boosts and is orthogonal to our contributions in this paper, left to be investigated, including quantization, model pruning, on-device optimization, mixed-precision inference, etc.. Besides, we haven’t applied any augmentation during training, thus, proper data or prompt augmentation might also be promising directions. We hope EdgeSAM will encourage more real-world applications with its on-device interactive segmentation capability and will continue working on further improvements.
6 Data Availability Statements
The datasets used in this study are publicly available as follows: